What is the positive solution of:
x^2 + 5x = 36 ?

Thank you so much in advance!

Answer by nerdybill(5411)

x^2 + 5x = 36
x^2 + 5x - 36 = 0
Factoring:
(x+9)(x-4) = 0
x = {-9,5}
.
The "positive" solution is then:
x = 5